# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 6.1 Klein four group case
Keep the notation as above. Assume that the hyperfocal subgroup $D$ is a Klein four group. Then the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to either $\mbox{$\mathcal{O}$}\mathrm{A}_{4}$ or the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. In particular, the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}(D\rtimes C_{3})$, where $C_{3}$ denotes a cyclic group of order $3$.
@@
For any nontrivial subgroup $Q$ of $D$, it is clear that the quotient group $D/Q$ is either trivial or a cyclic group of order $2$. In each case, the primitive Dade $D$-interior algebra $S(Q)$ is just $k$. Then the primitive Dade $D$-interior algebra $S$ has to be $k$ and the primitive $L$-interior algebra $\mathbb{A}^{\prime}$ in the paragraph 5.12 is just $\mathbb{A}_{\delta}$. By applying Theorem 5.14, we get the stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$ is just given by the restriction. By Corollary 4.7, there is at least one of the simple $\bar{\mathbb{A}}_{\delta}$-modules has odd dimension. Now we can apply the same arguments of the proof of [17, Proposition 12.2.9] to our setting and obtain that exactly one of the following statments holds.
@@
In the case (i), we have $\bar{\mathbb{A}}_{\delta}\mathop{\otimes}\limits_{kL}T_{\mathrm{i}}\cong\Omega^{-n}_{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{\mathrm{i}})$ for any $1\leq\mathrm{i}\leq 3$. Therefore by [16, Theorem 4.14.10], the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Hence, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual $M^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, here $j\in\delta$ (see the paragraph 3.1).
In the case (ii), we denote by $B_{0}(\mathrm{A}_{5})$ the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. The same argument in the last paragraph of the proof of [17, Theorem 12.1.2] shows that there is an indecomposable summand $N^{\prime}$ of the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ such that it and its dual induce a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$ and the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}B_{0}(\mathrm{A}_{5})^{\circ}}^{n}(N^{\prime})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$. Hence, the $\mathbb{A}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$ and its dual $N^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $B_{0}(\mathrm{A}_{5})$.
Finally, by [17, Theorem 12.4.1], $\mbox{$\mathcal{O}$}L$ and $B_{0}(\mathrm{A}_{5})$ are Rickard equivalent and then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}L$, which is just $\mbox{$\mathcal{O}$}(D\rtimes C_{3})$. ∎
6.5.
@@
By the structure theory of source algebras with normal defect group, the source algebra of the block $b_{0}$ is isomorphic to $\mbox{$\mathcal{O}$}(L\times R)$. By [7, Theorem 1.1 (i)], we have $l(b):=|\mathrm{IBr}(b)|=3=|E_{\mathfrak{h}}|$. Then by [31, Theorem], there is an isotypy between the blocks $b$ and $c$. Recall that $A$ and $\mathbb{A}$ are a source algebra and a hyperfocal subalgebra of the block $b$, respectively. Then by [9, Proposition 4.8], there is an $\mathcal{O}$-algebra isomorphism $A\cong\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$. It is clear that as $\mathcal{O}$-algebras, $\mbox{$\mathcal{O}$}(\mathrm{A}_{4}\times R)$ is isomorphic to $\mbox{$\mathcal{O}$}\mathrm{A}_{4}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$ and the principal block algebra of $\mbox{$\mathcal{O}$}(\mathrm{A}_{5}\times R)$ is isomorphic to $B_{0}(\mathrm{A}_{5})\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$, where $B_{0}(\mathrm{A}_{5})$ denotes the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. Then the statements of the proposition follows from Theorem 6.4. ∎
@@
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then we are in the case (i) of Theorem 6.4. In particular, the dimension of every simple $\bar{\mathbb{A}}_{\delta}$-module $\bar{W}_{\mathrm{i}}$ is coprime to $2$ by [17, Theorem 7.2.10]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=\sum\limits_{\mathrm{i}=1}^{3}[\bar{W}_{\mathrm{i}}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{1,2,3\}$. By the definition of the decomposition map $d_{\mathbb{A}_{\delta}}$, we have $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{4})=\sum\limits_{\mathrm{i}=1}^{3}\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ and $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})=\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ for for any $\mathrm{i}\in\{1,2,3\}$. In particular, every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case.
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. Then we are in the case (ii) of Theorem 6.4. In particular, $\mathrm{dim}_{k}(\bar{W}_{1})$ is coprime to $2$ and $\mathrm{dim}_{k}(\bar{W}_{2})=\mathrm{dim}_{k}(\bar{W}_{3})=2$ by [17, Corollary 7.2.11]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, of which the decomposition matrix is well-known, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{1}])=[\bar{W}_{1}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=[\bar{W}_{1}]+[\bar{W}_{2}]+[\bar{W}_{3}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{1}]+[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{2,3\}$. Hence, we still obtain that every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case. We are done. ∎
@@
Let $U$ be a finitely generated $\mathbb{A}_{\delta}$-module which is free as an $\mathcal{O}$-module. Then we have $\mathrm{rank}_{\mathcal{O}}(U)\equiv\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)~(\mathrm{mod}~2)$. In particular, $\mathrm{rank}_{\mathcal{O}}(U)$ is coprime to $2$ if and only if $\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ is coprime to $2$.
@@
First, note that $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is $\mathcal{O}$-free when $U$ is $\mathcal{O}$-free. So $\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ makes sense. Denote by $m_{\delta}$ the multiplicity of $\mathbb{A}$ at $\delta$ (see [19, Definition 2.2]). Then by [25, Propositions 3.3 and 3.5], we have $m_{\delta}$ coprime to $2$. Let $j^{\prime}$ be another element of $\delta$. Obviously, we have $j^{\prime}\mathbb{A}j$ isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}D$-$\mathbb{A}_{\delta}$-bimodules. Since $D_{\delta}$ is the unique local pointed group of $D$ on $\mathbb{A}$, there is an $\mbox{$\mathcal{O}$}D$-module decomposition
$\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U\cong U^{m_{\delta}}\oplus(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ (6.8.1)
@@
for some idempotent $e$ of $\mathbb{A}^{D}$ with $\mathrm{Br}_{D}(e)=0$. Here, $U^{m_{\delta}}$ denotes a direct sum of $m_{\delta}$-copies of $U$. Suppose that $e$ is primitive. By Rosenberg’s lemma, we have $e\in\mathrm{Tr}_{Q}^{D}(\mathbb{A}^{Q})$ for some proper subgroup $Q$ of $D$. By [29, Theorem 23.1], there is a primitive idempotent $e^{\prime}$ of $\mathbb{A}^{Q}$ such that $e=\mathrm{Tr}_{Q}^{D}(e^{\prime})$ and $e^{\prime}\cdot(e^{\prime})^{v}=(e^{\prime})^{v}\cdot e^{\prime}=0$ for any $v\in D-Q$. For any $\mathfrak{a}\in e\mathbb{A}j$ and any $x\in U$, the map sending $\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x$ to $\sum\limits_{uD\in D/Q}u\mathop{\otimes}\limits_{\mathcal{O}Q}(e^{\prime}u^{-1}\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x)$ determines an isomorphism of $\mbox{$\mathcal{O}$}D$-modules $e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}W_{\delta}\cong\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}(e^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$. In particular, $\mathrm{rank}_{\mathcal{O}}(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ is divided by $2$. In general, by decomposing $e$ as a sum of pairwise orthogonal primitive idempotents of $\mathbb{A}^{D}$, we still have $\mathrm{rank}_{\mathcal{O}}(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}W_{\delta})$ divided by $2$. Hence, by the decomposition (6.8.1), we have $m_{\delta}\mathrm{rank}_{\mathcal{O}}(U)\equiv\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)~(\mathrm{mod}~2)$. Then the lemma follows from $m_{\delta}$ coprime to $2$. ∎
@@
Combining Lemmas 6.7 and 6.8 with the Morita equivalence between $\mathbb{A}$ and $\mathbb{A}_{\delta}$, we can get the following result about the dimensions of simple $\hat{\mathbb{A}}-$modules.
@@
6.10.
Set $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{\hat{U}_{1},\hat{U}_{2},\hat{U}_{3},\hat{U}_{4}\}$ and $\mathrm{Irr}_{k}(\mathbb{A})=\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ if no confusion arises. By the proof of Lemma 6.7, we can describe the decomposition map $d_{\mathbb{A}}$ of $\mathbb{A}$ as follows. When $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, then for any $\mathrm{i}\in\{1,2,3\}$,
@@
for any $\mathrm{i}\in\{2,3\}$. It is clear that the conjugation action of $P$ on $\mathbb{A}$ can induce actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{k}(\mathbb{A})$, respectively. First, we demonstrate that $C_{P}(D)$ acts trivially on these two sets. This follows from the equivariant property of the Morita equivalences occurring in Theorem 6.4. For simplicity, we denote $C_{P}(D)$ by $P_{0}$.
Lemma 6.11.
Suppose that $P_{0}$ acts trivially on $L$ and $\mathrm{A}_{5}$. Then the Morita equivalences occurring in Theorem 6.4 are both $\Delta(P_{0})$-equivariant.
@@
We will borrow the notation from the proof of Theorem 6.4. Fix an element $u$ of $P_{0}$.

### v2

@@ 6.1 Klein four group case
Keep the notation as above. Assume that the hyperfocal subgroup $D$ is a Klein four group. Then the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to either $\mbox{$\mathcal{O}$}\mathrm{A}_{4}$ or the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. In particular, the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}\mathrm{A}_{4}$.
@@
For any nontrivial subgroup $Q$ of $D$, it is clear that the quotient group $D/Q$ is either trivial or a cyclic group of order $2$. In each case, the primitive Dade $D$-interior algebra $S(Q)$ in the paragraph 5.6 is just $k$. Then the primitive Dade $D$-interior algebra $S$ has to be the trivial $D$-interior algebra $\mathcal{O}$ and the primitive $L$-interior algebra $\mathbb{A}^{\prime}$ in the paragraph 5.12 is just $\mathbb{A}_{\delta}$. By applying Theorem 5.14, we get that a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$ which is induced by the restriction. By Corollary 4.7, there is at least one of the simple $\bar{\mathbb{A}}_{\delta}$-modules of odd dimension. Therefore, the proof of [16, Proposition 12.2.9] still applies to to our setting and then exactly one of the following statments holds.
@@
In the case (i), we have $\bar{\mathbb{A}}_{\delta}\mathop{\otimes}\limits_{kL}T_{\mathrm{i}}\cong\Omega^{-n}_{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{\mathrm{i}})$ for any $1\leq\mathrm{i}\leq 3$. Therefore by [15, Theorem 4.14.10], the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Hence, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual $M^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$.
In the case (ii), we denote by $B_{0}(\mathrm{A}_{5})$ the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. Since $L$ is isomorphic to ${\rm A}_{4}$, we can view $L$ as a subgroup of ${\rm A}_{5}$. The same argument in the last paragraph of the proof of [16, Theorem 12.1.2] shows that there is an indecomposable summand $N^{\prime}$ of the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ such that it and its dual induce a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$ and the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}B_{0}(\mathrm{A}_{5})^{\circ}}^{n}(N^{\prime})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$. Hence, the $\mathbb{A}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$ and its dual $N^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $B_{0}(\mathrm{A}_{5})$.
Finally, by [16, Theorem 12.4.1], $\mbox{$\mathcal{O}$}L$ and $B_{0}(\mathrm{A}_{5})$ are Rickard equivalent and then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}L$, which is just $\mbox{$\mathcal{O}$}\mathrm{A}_{4}$. ∎
$6.$
@@
By the structure theory of source algebras with normal defect groups, the source algebra of the block $b_{0}$ is isomorphic to $\mbox{$\mathcal{O}$}(L\times R)$ as $P$-interior algebras. Denote $|\mathrm{IBr}(b)|$ by $l(b)$. By [7, Theorem 1.1 (i)], we have $l(b)=3=|E_{\mathfrak{h}}|$. Then by [31, Theorem], there is an isotypy between the blocks $b$ and $b_{0}$. Recall that $A$ and $\mathbb{A}$ are a source algebra and a hyperfocal subalgebra of the block $b$, respectively. Then by [9, Proposition 4.8], there is an $\mathcal{O}$-algebra isomorphism $A\cong\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$. It is clear that as $\mathcal{O}$-algebras, $\mbox{$\mathcal{O}$}(\mathrm{A}_{4}\times R)$ is isomorphic to $\mbox{$\mathcal{O}$}\mathrm{A}_{4}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$ and the principal block algebra of $\mbox{$\mathcal{O}$}(\mathrm{A}_{5}\times R)$ is isomorphic to $B_{0}(\mathrm{A}_{5})\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$. Then the statements of the proposition follow from Theorem 6.5. ∎
@@
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then we are in the case (i) of Theorem 6.5. In particular, the dimension of every simple $\bar{\mathbb{A}}_{\delta}$-module $\bar{W}_{\mathrm{i}}$ is coprime to $2$ by [16, Theorems 7.2.10 (ii) and 7.2.1 (ii)]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=\sum\limits_{\mathrm{i}=1}^{3}[\bar{W}_{\mathrm{i}}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{1,2,3\}$. By the definition of the decomposition map $d_{\mathbb{A}_{\delta}}$, we have $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{4})=\sum\limits_{\mathrm{i}=1}^{3}\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ and $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})=\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ for any $\mathrm{i}\in\{1,2,3\}$. In particular, every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case.
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. Then we are in the case (ii) of Theorem 6.5. In particular, $\mathrm{dim}_{k}(\bar{W}_{1})$ is coprime to $2$ and $\mathrm{dim}_{k}(\bar{W}_{2})=\mathrm{dim}_{k}(\bar{W}_{3})=2$ by [16, Corollary 7.2.11]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, of which the decomposition matrix is well-known, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{1}])=[\bar{W}_{1}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=[\bar{W}_{1}]+[\bar{W}_{2}]+[\bar{W}_{3}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{1}]+[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{2,3\}$. Hence, we still obtain that every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case. We are done. ∎
@@
Let $U$ be a finitely generated $\mathbb{A}_{\delta}$-module which is free as an $\mathcal{O}$-module. Then the $\mathcal{O}$-rank of $U$ is coprime to $2$ if and only if the $\mathcal{O}$-rank of $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is coprime to $2$.
@@
First, note that $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is $\mathcal{O}$-free since the $\mathbb{A}$-$\mathbb{A}_{\delta}$-bimodule $\mathbb{A}j$ induces a Morita equivalence between $\mathbb{A}$ and $\mathbb{A}_{\delta}$. So the notation $\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ makes sense. Denote by $m_{\delta}$ the multiplicity of $\mathbb{A}$ at $\delta$ (see [19, Definition 2.2]). Then by [25, Propositions 3.3 and 3.5], we have $m_{\delta}$ coprime to $2$. Let $j^{\prime}$ be another element of $\delta$. Obviously, we have $j^{\prime}\mathbb{A}j$ isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}D$-$\mathbb{A}_{\delta}$-bimodules. Since $D_{\delta}$ is the unique local pointed group of $D$ on $\mathbb{A}$, there is an $\mbox{$\mathcal{O}$}D$-module decomposition
$\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U\cong U^{m_{\delta}}\oplus({\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ (6.9.1)
@@
for some idempotent $\rm e$} of $\mathbb{A}^{D}$ with $\mathrm{Br}_{D}({\rm e})=0$. Here, $U^{m_{\delta}}$ denotes a direct sum of $m_{\delta}$-copies of $U$. Suppose that $\rm e$} is primitive. By Rosenberg’s lemma, we have ${\rm e}\in\mathrm{Tr}_{Q}^{D}(\mathbb{A}^{Q})$ for some proper subgroup $Q$ of $D$. By [29, Theorem (23.1)], there is a primitive idempotent $\rm e$$^{\prime}$ of $\mathbb{A}^{Q}$ such that ${\rm e}=\mathrm{Tr}_{Q}^{D}({\rm e}^{\prime})$ and ${\rm e}^{\prime}\cdot({\rm e}^{\prime})^{v}=({\rm e}^{\prime})^{v}\cdot{\rm e}^{\prime}=0$ for any $v\in D-Q$. For any $\mathfrak{a}\in{\rm e}\mathbb{A}j$ and any $x\in U$, the map sending $\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x$ to $\sum\limits_{uQ\in D/Q}u\mathop{\otimes}\limits_{\mathcal{O}Q}({\rm e}^{\prime}u^{-1}\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x)$ determines an isomorphism of $\mbox{$\mathcal{O}$}D$-modules ${\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U\cong\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}({\rm e}^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$. Since ${\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is isomorphic to a direct summand of $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ as $\mbox{$\mathcal{O}$}D$-modules, ${\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is also a free $\mathcal{O}$-module of finite rank. Furthermore, it is clear that $\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}({\rm e}^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ is a direct sum of $|D/Q|$-copies of ${\rm e}^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ as $\mathcal{O}$-modules. Hence, ${\rm e}^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is a free $\mathcal{O}$-module of finite rank and then $\mathrm{rank}_{\mathcal{O}}({\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)=|D/Q|\cdot\mathrm{rank}_{\mathcal{O}}({\rm e}^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$, which is divided by $2$ since $Q$ is a proper subgroup of $D$.
In general, by decomposing $\rm e$} as a sum of pairwise orthogonal primitive idempotents of $\mathbb{A}^{D}$, we still have $\mathrm{rank}_{\mathcal{O}}({\rm e}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ divided by $2$. Hence, by the decomposition (6.9.1), we have $m_{\delta}\mathrm{rank}_{\mathcal{O}}(U)\equiv\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)~(\mathrm{mod}~2)$. Then the lemma follows from $m_{\delta}$ being coprime to $2$. ∎
Remark 6.10.
In fact, Lemma 6.9 is true for any prime $p$ and any hyperfocal subgroup $D$ since the proof above does not depend on the assumption that $D$ is a Klein four group.
@@
Combining Lemmas 6.8 and 6.9 with the Morita equivalence between $\mathbb{A}$ and $\mathbb{A}_{\delta}$, we can get the following result about the dimensions of simple $\hat{\mathbb{A}}-$modules.
@@
6.12.
Set $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{\hat{U}_{1},\hat{U}_{2},\hat{U}_{3},\hat{U}_{4}\}$ and $\mathrm{Irr}_{k}(\mathbb{A})=\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ if no confusion arises. By the proof of Lemma 6.8, we can describe the decomposition map $d_{\mathbb{A}}$ of $\mathbb{A}$ as follows. When $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}{\rm A}_{4}$, then for any $\mathrm{i}\in\{1,2,3\}$,
@@
for any $\mathrm{i}\in\{2,3\}$. It is clear that the conjugation action of $P$ on $\mathbb{A}$ can induce the conjugation actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{k}(\mathbb{A})$, respectively.
6.13.
On the other hand, since $L$ is isomorphic to ${\rm A}_{4}$, we can view $L$ as a subgroup of ${\rm A}_{5}$ such that $D$ equals $\{(1),(12)(34),(13)(24),(14)(23)\}$ and $E_{\mathfrak{h}}$ equals $\{(1),(123),(132)\}$. We denote $C_{P}(D)$ by $P_{0}$. It is clear that $P/P_{0}$ is isomorphic to a $2$-subgroup of the automorphism group ${\rm Aut}(D)$ of $D$. Then $P/P_{0}$ is either trivial or a cyclic group of order $2$. We can define an action of $P/P_{0}$ on ${\rm A}_{5}$ which stabilizes $L$ as follows. When $P/P_{0}$ has order $2$, the action of $P/P_{0}$ on ${\rm A}_{5}$ is given by the conjugation action of the transposition $(12)$ in ${\rm S}_{5}$. Here, ${\rm S}_{5}$ denotes the symmetric group on $5$ letters. So it can induce an action of $P$ on ${\rm A}_{5}$ through the canonical map $P\longrightarrow P/P_{0}$ such that this action stabilizes $L$ and the principal block $B_{0}({\rm A}_{5})$. Moreover, set $P/P_{0}=\langle u_{0}P_{0}\rangle$ when $P/P_{0}$ is nontrivial. Then $T_{1}$ is the unique $P$-stable kLkL-simple module. $T_{2}$ and $T_{3}$ are both $P_{0}$-stable $kL$-simple modules and ${{}_{u_{0}}}T_{2}=T_{3}$. Note that in both cases, the actions of $P/P_{0}$ on $L=D\rtimes E_{\mathfrak{h}}$ are the same as the conjugation actions of $P/P_{0}$ on $L$. Now we can get the following $\Delta(P)$-equivariant property of the Morita equivalences occurring in Theorem 6.5.
Proposition 6.14.
Keep the notation as above. Then the Morita equivalences occurring in Theorem 6.5 are both $\Delta(P)$-equivariant, namely that the bimodules induce the Morita equivalences occurring in Theorem 6.5 are $\Delta(P)$-equivariant.
@@
We will borrow the notation from the paragraph 4.9 and the proof of Theorem 6.5 and Lemma 5.18. Fix an element $u$ of $P$. By Lemma 5.18, there is an element $a_{u,\delta}$ of $A_{\delta}^{\times}$ such that ${{}^{a_{u,\delta}}}\mathfrak{e}_{\tilde{g}}=\mathfrak{e}_{{{}^{u}}\tilde{g}}$ for any $\tilde{g}\in E_{\mathfrak{h}}$.